A 200 g, 40 cm-diameter turntable rotates on frictionless bearings at 60 rpm (rounds per minute). A 10 g block sits at the center of the turntable. A compressed spring shoots the block radially outward along a frictionless groove in the surface of the turntable. You may treat the block as a point mass. Calculate the rotational energy before and after the spring release. Is energy conserved?
Answer #1

Moment of inertia of turntable

It =(1/2)MtR2 =(1/2)*0.2*(0.4/2)2=4*10-3 Kg-m2

Moment of inertia of block

IB=MBR2=(0.01)(0.4/2)2=4*10-4 Kg-m2

By Conservation of angular Momentum

I1W1=I2W2

(4*10-3)*60 =(4*10-4+4*10-3)W2

W2=54.55 rpm

in rad/s

W1=60*2pi/60 =6.283 rad/s

W2=54.55*2pi/60 =5.712 rad/s

Rotational Kinetic energy before and and after

Kbefore=(1/2)I1W12 =(1/2)*(4*10-3)*6.2832 =0.078957 J =0.079 J

Kafter= (1/2)I2W22=(1/2)*(4.4*10-3)*5.7122=0.071779 J =0.0718 J

NO energy is not conserved